Problem 1 The percent of fat calories that a person in America consumes each day is normally distributed with a mean of about 32 and a standard deviation of about 12 Suppose that 18 individuals are randomly chosen. Let X = average percent of fat calories consumed for a sample of size 18 Enter numbers as integers or fractions in p/q form, or as decimals accurate to nearest 0.01 a.(20) X ~ B ( 32 , 12v18 ) . b. (20) Use the mean and standard deviation of X to determine the z value for X 35 c. (20) mustrate PX < 35) as an area by adjusting the slider along the horizontal axis to control the z value. 0.00 0.300 0.2 0.1
z value. c. (20) Illustrate P(X < 35) as an area by adusting the slider along the horizontal axis to control the z- 0.00 Expand A 300 0.2 0.1 oill left Fill right d (.20) What i the probablitythat the average fat consumption for 18 people in America is less than 35 P(K <35) e (20) Find the trst quartile for the average percent of fat calories
Answer #1

Let X be the percent of fat calories that a person in America consumes each day. X is normally distributed with mean μ=32 and standard deviation 12

A sample of size n=18 individuals is randomly chosen. Let ar{X} be the average percent of fat calories consumed for a sample of size n=18.

We know that ar{X} is normally distributed with mean mu_{ar{x}}=mu=32 and standard deviation (or also called standard error)  

3

a) ans: X N (32,12/V18) fractions in"p/q" format.

or in decimals as

X ~N(32,2.83)

b) The z value for X=35 is

35- 35-32 2.83 1 .06

c) The value of z is 1.06. We need to move the slider along to the right till we get to z=1.06 as shown below. The shaded region to the left of the vertical line would be the region corresponding to X 〈 35 or Z<1.06

0.00 0.300 に1.06 0.2 0.1 OFill left Fll right

d) The probability that the average consumption for 18 people in America is less than 35 is

P(X < 35) = P(Z < 1.06) We represent 35 using its z value

Using the standard normal table we read the value for z=1.06 as 0.3554

The value of P(Z<1.06) = 0.5+0.3553=0.8553

ans: egin{align*} P(ar{X}<35)=0.86 end{align*}

e) Let egin{align*} Q_1 end{align*} be the first quartile for the average percent of fat calories.

We know that 25% of the observations are less than the first quartile egin{align*} Q_1 end{align*} . That is the probability that the average consumption for 18 people in America is less than egin{align*} Q_1 end{align*} is 0.25

P(X < Q1) 0.25

First we will find the z value for which P(Z<z) = 0.25. We know that the probability that Z is less than 0 is 0.5. That is P(Z<0) = 0.5

Hence z has to be less than zero (negative) for the probability to be less than 0.5.

Hence we will write the required probability as

egin{align*} &P(Z<-z)=0.25 implies &P(Z>z)=0.25quad ext{this is due to the symmetry of normal distribution} implies &1-P(Z<z)=0.25 implies &P(Z<z)=1-0.25 implies &P(Z<z)=0.75 end{align*}

Now we will use the standard normal table and find that for z=0.67 the area under the normal curve is (0.5+0.2486)=0.7486. Since this is close enough to 0.75, we will use this value. (For more accurate value, for z=0.68 the area under the normal curve is 0.5+0.2517=0.7517. Now we interpolate and get a more accurate 0.674)

That is P(Z<0.67) = 0.75 or P(Z<-0.67) = 0.25

Now we will get the value of egin{align*} Q_1 end{align*} by equating the z value of egin{align*} Q_1 end{align*} to -0.67

1-32 2.83 is the z value, which we will equate to -0.67

egin{align*}&rac{Q_1-32}{2.83}=-0.67 implies &Q_1=32-2.83 imes 0.67 implies &Q_1=30.10 end{align*}

(If we use a more accurate value of 0.674 we will get the value of Q1=30.09)